need help with these two rxns please
NaNO2, HCl (CH3CH2)2NH, HT 2 CH3 CH2CH3
Answer #1

NH2 NONO 2. HLE BENE I + H2O 0°S U13 CH3 Diazonium salt Methanism U NaNO2 + Hue - H-O-N=0 i H-O-N =ö + H@ — Hoyo AS NE860 (iiPark li) H2 CH3 NHT CHg (4.3 995 an Enamine formation Reaten belauge is a bewnd any amine Mechanism (الماد { + N 544 2043 d